# From Image Denoisers to Regularizing Imaging Inverse Problems: An Overview

### v1

@@ Abstract
Inverse problems lie at the heart of modern imaging science, with broad applications in areas such as medical imaging, remote sensing, and microscopy. Recent years have witnessed a paradigm shift in solving imaging inverse problems, where data-driven regularizers are used increasingly, leading to remarkably high-fidelity reconstruction. A particularly notable approach for data-driven regularization is to use learned image denoisers as implicit priors in iterative image reconstruction algorithms. This survey presents a comprehensive overview of this powerful and emerging class of algorithms, commonly referred to as plug-and-play (PnP) methods. We begin by providing a brief background on image denoising and inverse problems, followed by a short review of traditional regularization strategies. We then explore how proximal splitting algorithms, such as the alternating direction method of multipliers (ADMM) and proximal gradient descent (PGD), can naturally accommodate learned denoisers in place of proximal operators, and under what conditions such replacements preserve convergence. The role of Tweedie’s formula in connecting optimal Gaussian denoisers and score estimation is discussed, which lays the foundation for regularization-by-denoising (RED) and more recent diffusion-based posterior sampling methods. We discuss theoretical advances regarding the convergence of PnP algorithms, both within the RED and proximal settings, emphasizing the structural assumptions that the denoiser must satisfy for convergence, such as non-expansiveness, Lipschitz continuity, and local homogeneity. We also address practical considerations in algorithm design, including choices of denoiser architecture and acceleration strategies. By integrating both classical optimization insights and modern learning-based priors, this survey aims to provide a unified and accessible framework for understanding PnP methods and their theoretical underpinnings.
@@ 1 Introduction
Thanks to the pioneering work by Venkatakrishnan et al. venkat_pnp_6737048 , plug-and-play (PnP) methods have emerged in the last decade as an effective paradigm that deviates from handcrafted priors by integrating powerful image denoisers directly into iterative optimization algorithms for solving inverse problems (see kamilov2023plug for a recent survey on PnP methods). The key insight behind PnP is deceptively simple yet quite powerful: instead of specifying an explicit prior, one can “plug in” an off-the-shelf image denoiser within an iterative framework such as alternating direction method of multipliers (ADMM) or proximal gradient descent (PGD), effectively regularizing the solution through repeated denoising operations. This idea has led to a family of flexible algorithms, with PnP-ADMM chan2016plug and PnP-PGD pmlr-v97-ryu19a ; convergent_prox_pnp_pgd being two of the most prominent and widely studied variants of PnP algorithms. PnP methods are particularly attractive due to their modularity; the same image denoiser leads to a reasonable image reconstruction for different forward operators. By decoupling the forward model from the prior, PnP enables practitioners to exploit state-of-the-art denoisers, whether model-based (such as BM3D bm3d_main ) or data-driven (e.g. deep convolutional neural network (CNN)-based denoisers dncnn_main ; pnp_deep_denoiser ), without the need to derive complex closed-form priors or incorporate explicit regularizers into the variational framework. This practical plug-and-play nature makes these methods highly attractive and flexible for diverse applications where noise statistics, measurement models, and image features vary widely.
@@
In parallel, the regularization by denoising (RED) approach  red_romano ; red_schniter offers another compelling framework by constructing a class of explicit regularizers using off-the-shelf image denoisers. RED provides a unifying view that connects variational principles with denoising-based operators, offering fixed-point iterations with interpretability and, under suitable conditions, convergence guarantees red_schniter . Other notable theoretical advances include rigorous convergence guarantees for PnP-PGD and PnP-ADMM under appropriate Lipschitz continuity assumptions on the denoising residual (see Theorems 1, 2, and Corollary 3 in pmlr-v97-ryu19a ), and the extension of PnP concepts to broader classes of iterative schemes, such as consensus equilibrium buzzard2018plug and block coordinate methods red_block_coord . The synergy between PnP frameworks and deep learning has further amplified their impact. Learned denoisers, such as DnCNN dncnn_main , FFDNet ffdnet , and more recent transformer-based pnp_transformer or denoising diffusion model-based approaches pnp_denoising_diffusion can be seamlessly integrated into PnP schemes, delivering remarkable reconstruction performance even in severely ill-posed settings. Beyond static images, PnP methods are increasingly being adapted to dynamic and multimodal imaging tasks pnp_dynamic_NLOS , where additional temporal or cross-modality constraints may be incorporated, while still retaining the modular plug-and-play philosophy.
@@
Despite their empirical success, PnP methods still present open challenges. Key questions include understanding their theoretical guarantees when using highly nonlinear or non-expansive denoisers, designing adaptive schemes that can select or learn denoising strength on the fly, and extending the framework to handle non-Gaussian noise, physics-based constraints, or multimodal data fusion. In this survey, we aim to provide a comprehensive and up-to-date account of plug-and-play methods for imaging inverse problems. We trace the historical development of this framework, highlight the underlying theoretical foundations, and discuss advances in algorithmic design and learning-based denoisers. We also discuss the application of PnP denoisers for Bayesian imaging that incorporates denoisers with diffusion-based posterior sampling, and identify open research directions that may shape the future of PnP-based inverse problem solving. We focus particularly on methods that are widely regarded as pioneering in the area of PnP imaging and methods that come with rigorous convergence guarantees.
@@ 1.1 Brief Survey of Image Denoising
Image denoising has long been a fundamental problem in signal and image processing, driving the development of increasingly sophisticated models and algorithms over the past several decades. We refer interested readers to GOYAL2020220_denoise_elsv ; elad_denoising_review ; is_denoising_dead for a review on image denoisers, covering classical techniques to modern deep learning-based approaches. In the discussion that follows, we will assume that the image $x$ is discretized, and can therefore be represented as a vector in $\mathbb{R}^{n}$, assuming that the image has $n$ pixels. Under an additive noise model, the goal of image denoising is to recover an unknown clean image $x\in\mathbb{R}^{n}$ from a noisy observation $z=x+w$, where $w$ represents additive noise. Depending on the application, $w$ may follow a Gaussian or non-Gaussian distribution, and may be white or colored. The fundamental challenge arises from the fact that noise corrupts both low- and high-frequency components in the image, making it difficult to separate noise from fine image details. Designing an effective denoising algorithm therefore requires balancing noise suppression with the preservation of edges, textures, and small structures in the underlying image. An important feature of well-designed (though not necessarily perfect) denoisers is that they can naturally generate a multiscale decomposition of an image Milanfar2025-xp , while still allowing exact reconstruction. This idea parallels classical multiscale decompositions, such as the Laplacian pyramid laplace_pyramid , but is achieved here through the action of the denoiser.
@@
$\left(B,\left(\boldsymbol{\alpha}_{i}\right)_{i=1}^{p}\right)=\arg\min_{B,\left(\boldsymbol{\alpha}_{i}\right)_{i=1}^{p}}\sum_{i=1}^{p}\left[\|z_{i}-B\boldsymbol{\alpha}_{i}\|_{2}^{2}+\lambda_{i}\|\boldsymbol{\alpha}_{i}\|_{1}\right],$
@@
Throughout these decades of progress, one recurring theme has been the interplay between model-based image priors and data-driven learning of denoisers. Classical algorithms offer interpretability, well-defined mathematical properties, and provable convergence guarantees, but often lack the representational power needed to capture the complexity of natural images. Learned denoisers excel at modeling rich high-dimensional distributions but raise questions about stability, generalization, and robustness under distribution shifts. This tension has inspired the design of hybrid approaches: These methods embed powerful learned denoisers into model-based optimization frameworks. This idea gave rise to the family of plug-and-play (PnP) methods, which have become highly successful because they bring together two advantages: the flexibility and performance of learned denoisers, and the interpretability and control offered by classical iterative schemes.
@@ 1.2 Inverse Problems and Regularization
where $K:\mathcal{X}\rightarrow\mathcal{Y}$ is a compact operator between Hilbert spaces, and $w$ represents measurement noise. The ill-posed nature manifests through three distinct aspects of the operator-theoretic framework: either injectivity or surjectivity of the forward operator may not hold, or stability of the solution map might be violated. For instance, if $K$ is a compact operator with an infinite-dimensional range, then surjectivity and stability are violated. This is, for instance, the case for the ray transform operator that underlies many medical imaging modalities, such as computed tomography (CT) and positron emission tomography (PET). To see this, first consider the case where the range $\mathscr{R}(K)$ of the forward operator $K$ is not closed, implying that solutions may not exist for an arbitrary $y\in\mathcal{Y}$, as any non-zero noise component $w$ orthogonal to $\mathscr{R}(K)$ renders the problem unsolvable in the strict sense. Second, the potential non-triviality of $\mathscr{N}(K)$, the null-space of the forward operator $K$, violates uniqueness, particularly evident in limited-angle tomography where certain features become invisible. Most critically, the unboundedness of the generalized inverse $A^{\dagger}$, when restricted to $\mathscr{R}(K)$, leads to extreme sensitivity to noise, leading to an unstable solution (i.e., a small amount of noise in the measurement yields a drastically different solution).
@@
where the functional $R:\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ encodes prior knowledge about plausible solutions. The classical Tikhonov regularization tikhonov_main employs $R(x)=\|x\|_{\mathcal{X}}^{2}$, yielding the explicit solution $x_{\alpha}=(K^{*}K+\alpha I)^{-1}K^{*}y$, which corresponds to a spectral filter with coefficients $f_{\alpha}(\sigma)=\sigma/(\sigma^{2}+\alpha)$. While the quadratic ($L^{2}$) penalty on $x$ guarantees stability of the reconstruction, it does so by uniformly penalizing deviations across all frequencies. This effect can be seen from the Euler–Lagrange optimality condition $K^{*}(Kx-y)+\alpha x=0$, where the regularization parameter $\alpha$ acts as a frequency-independent damping term. The outcome is a systematic suppression of both noise and fine-scale features, resulting in overly smoothed reconstructions and noticeable loss of sharp edge structures. To mitigate this limitation, one may instead consider sparsity-promoting penalties such as the $L^{1}$ norm lasso_tibshirani . Unlike the quadratic case, the $L^{1}$ norm is non-differentiable at zero, which effectively allows the reconstruction to tolerate large local variations and thus preserve sharp discontinuities such as edges. However, this comes with new drawbacks: the piecewise-constant bias induced by $L^{1}$ regularization can lead to the well-known staircasing artifact, where smooth gradients in the original image are replaced by artificial flat regions separated by sharp transitions of intensity RUDIN1992259_tv ; Chambolle1997_tv_related . Moreover, textures and fine oscillatory details tend to be lost, since they are not easily represented in the sparse model enforced by the $L^{1}$ norm.
@@ 1.3 Convergent Regularization
To obtain stable solutions to inverse problems, a mechanism is needed to handle varying noise levels in the measurement. For this purpose, the concept of convergent regularization has proven highly useful. Regularization can be roughly understood as a convergence requirement to a unique solution, such as the minimum-norm solution $x^{\dagger}$, where convergence occurs as the noise level $\delta\rightarrow 0$. Formally, consider the previously discussed reconstruction operator $\mathcal{R}_{\alpha}\coloneqq\mathcal{R}(\cdot,\alpha)$, which parametrizes a family of continuous operators $\mathcal{R}_{\alpha}:\mathcal{Y}\to\mathcal{X}$. The parameter $\alpha$ depends on the noise level $\delta>0$, where $\|y^{\delta}-y^{0}\|\leq\delta$ and $y^{0}:=Kx^{*}$ denotes noise-free measurement data. We say that the family of reconstruction operators is a convergent regularization method if there exists a parameter choice rule $\alpha=\alpha(\delta,y^{\delta})$ such that reconstructions $x^{\delta}\coloneqq\mathcal{R}_{\alpha(\delta,y^{\delta})}(y^{\delta})$ converge to $x^{\dagger}\coloneqq A^{\dagger}y^{0}$ (given by the pseudo-inverse) as noise vanishes, in the sense that
@@
In other words, we have point-wise convergence of the reconstruction operators to the pseudo-inverse, i.e. $\mathcal{R}_{\alpha(\delta,y^{\delta})}(y^{\delta})\to A^{\dagger}y^{0}$ as $\delta\to 0$. We refer interested readers to engl1996regularization ; scherzer2009variational ; benning2018modern for a detailed discussion on convergent regularization schemes (and several convergence rate results in the classical regularization literature). While this is somewhat restrictive as it only considers convergence to the least-squares minimum-norm solution, this nevertheless serves as an important tool to design learned regularization methods, i.e., learned reconstruction approaches that formally satisfy the above convergence criterion.

### v2

# Solving Imaging Inverse Problems Using Plug-and-Play Denoisers: Regularization and Optimization Perspectives

@@ Abstract
Inverse problems lie at the heart of modern imaging science, with broad applications in areas such as medical imaging, remote sensing, and microscopy. Recent years have witnessed a paradigm shift in solving imaging inverse problems, where data-driven regularizers are used increasingly, leading to remarkably high-fidelity reconstruction. A particularly notable approach for data-driven regularization is to use learned image denoisers as implicit priors in iterative image reconstruction algorithms. This chapter presents a comprehensive overview of this powerful and emerging class of algorithms, commonly referred to as plug-and-play (PnP) methods. We begin by providing a brief background on image denoising and inverse problems, followed by a short review of traditional regularization strategies. We then explore how proximal splitting algorithms, such as the alternating direction method of multipliers (ADMM) and proximal gradient descent (PGD), can naturally accommodate learned denoisers in place of proximal operators, and under what conditions such replacements preserve convergence. The role of Tweedie’s formula in connecting optimal Gaussian denoisers and score estimation is discussed, which lays the foundation for regularization-by-denoising (RED) and more recent diffusion-based posterior sampling methods. We discuss theoretical advances regarding the convergence of PnP algorithms, both within the RED and proximal settings, emphasizing the structural assumptions that the denoiser must satisfy for convergence, such as non-expansiveness, Lipschitz continuity, and local homogeneity. We also address practical considerations in algorithm design, including choices of denoiser architecture and acceleration strategies. By integrating both classical optimization insights and modern learning-based priors, this chapter aims to provide a unified and accessible framework for understanding PnP methods and their theoretical underpinnings.
@@ 1 Introduction
Thanks to the pioneering work by Venkatakrishnan et al. venkat_pnp_6737048 , plug-and-play (PnP) methods have emerged in the last decade as an effective paradigm, deviating from handcrafted priors by instead integrating powerful image denoisers directly into iterative optimization algorithms for solving inverse problems (see kamilov2023plug for a recent survey on PnP methods). The key insight behind PnP is deceptively simple yet quite powerful: instead of specifying an explicit prior, one can “plug in” an off-the-shelf image denoiser within an iterative framework such as alternating direction method of multipliers (ADMM) or proximal gradient descent (PGD), effectively regularizing the solution through repeated denoising operations. This idea has led to a family of flexible algorithms, with PnP-ADMM chan2016plug and PnP-PGD pmlr-v97-ryu19a ; convergent_prox_pnp_pgd being two of the most prominent and widely studied variants of PnP algorithms. PnP methods are particularly attractive due to their modularity; the same image denoiser leads to a reasonable image reconstruction for different forward operators. By decoupling the forward model from the prior, PnP enables practitioners to exploit state-of-the-art denoisers, whether model-based (such as BM3D bm3d_main ) or data-driven (e.g. deep convolutional neural network (CNN)-based denoisers dncnn_main ; pnp_deep_denoiser ), without the need to derive complex closed-form priors or incorporate explicit regularizers into the variational framework. This practical plug-and-play nature makes these methods highly attractive and flexible for diverse applications where noise statistics, measurement models, and image features vary widely.
@@
In parallel, the regularization by denoising (RED) approach  romano2017little ; red_schniter offers another compelling framework by constructing a class of explicit regularizers using off-the-shelf image denoisers. RED provides a unifying view that connects variational principles with denoising-based operators, offering fixed-point iterations with interpretability and, under suitable conditions, convergence guarantees red_schniter . Other notable theoretical advances include rigorous convergence guarantees for PnP-PGD and PnP-ADMM under appropriate Lipschitz continuity assumptions on the denoising residual (see Theorems 1, 2, and Corollary 3 in pmlr-v97-ryu19a ), and the extension of PnP concepts to broader classes of iterative schemes, such as consensus equilibrium buzzard2018plug and block coordinate methods red_block_coord . The synergy between PnP frameworks and deep learning has further amplified their impact. Learned denoisers, such as DnCNN dncnn_main , FFDNet ffdnet , and more recent transformer-based pnp_transformer or denoising diffusion model-based approaches pnp_denoising_diffusion can be seamlessly integrated into PnP schemes, delivering remarkable reconstruction performance even in severely ill-posed settings. Beyond static images, PnP methods are increasingly being adapted to dynamic and multimodal imaging tasks pnp_dynamic_NLOS , where additional temporal or cross-modality constraints may be incorporated, while still retaining the modular plug-and-play philosophy.
@@
Despite their empirical success, PnP methods still present open challenges. Key questions include understanding their theoretical guarantees when using highly nonlinear or non-expansive denoisers, designing adaptive schemes that can select or learn denoising strength on the fly, and extending the framework to handle non-Gaussian noise, physics-based constraints, or multimodal data fusion. In this chapter, we aim to provide a comprehensive and up-to-date account of plug-and-play methods for imaging inverse problems. We trace the historical development of this framework, highlight the underlying theoretical foundations, and discuss advances in algorithmic design and learning-based denoisers. We also discuss the application of PnP denoisers for Bayesian imaging that incorporates denoisers with diffusion-based posterior sampling, and identify open research directions that may shape the future of PnP-based inverse problem solving. We focus particularly on methods that are widely regarded as pioneering in the area of PnP imaging and methods that come with rigorous convergence guarantees.
@@ 1.1 Brief Survey of Image Denoising
Image denoising has long been a fundamental problem in signal and image processing, driving the development of increasingly sophisticated models and algorithms over the past several decades. We refer interested readers to GOYAL2020220_denoise_elsv ; elad_denoising_review ; is_denoising_dead for a review on image denoisers, covering classical techniques to modern deep learning-based approaches. In the discussion that follows, we will assume that the image $x$ is discretized, and can therefore be represented as a vector in $\mathbb{R}^{n}$ after concatenating all the pixels and the color channels in the image. For a grayscale image of size $_{1}\times n_{2}, n=n$$n_{1}n_{2}$; and for a color image (with three color channels) of the same size, $n=3n_{1}n_{2}$. Under an additive noise model, the goal of image denoising is to recover an unknown clean image $x\in\mathbb{R}^{n}$ from a noisy observation $z=x+w$, where $w$ represents additive noise. Depending on the application, $w$ may follow a Gaussian or non-Gaussian distribution, and may be white or colored. The fundamental challenge arises from the fact that noise corrupts both low- and high-frequency components in the image, making it difficult to separate noise from fine image details. Designing an effective denoising algorithm therefore requires balancing noise suppression with the preservation of edges, textures, and small structures in the underlying image. An important feature of well-designed (though not necessarily perfect) denoisers is that they can naturally generate a multiscale decomposition of an image Milanfar2025-xp , while still allowing exact reconstruction. This idea parallels classical multiscale decompositions, such as the Laplacian pyramid laplace_pyramid , but is achieved here through the action of the denoiser.
@@
$\left(B,\left(\boldsymbol{\alpha}_{i}\right)_{i=1}^{p}\right)=\operatorname*{arg\,min}_{B,\left(\boldsymbol{\alpha}_{i}\right)_{i=1}^{p}}\sum_{i=1}^{p}\left[\|z_{i}-B\boldsymbol{\alpha}_{i}\|_{2}^{2}+\lambda_{i}\|\boldsymbol{\alpha}_{i}\|_{1}\right],$
@@
Throughout these decades of progress, one recurring theme has been the interplay between model-based image priors and data-driven learning of denoisers. Classical algorithms offer interpretability, well-defined mathematical properties, and provable convergence guarantees, but often lack the representational power needed to capture the complexity of natural images. Learned denoisers excel at modeling rich high-dimensional distributions but raise questions about stability, generalization, and robustness under distribution shifts. This tension has inspired the design of hybrid approaches, which embed powerful learned denoisers into model-based optimization frameworks. This idea gave rise to the family of plug-and-play (PnP) methods, which have become highly successful because they bring together two advantages: the flexibility and performance of learned denoisers, and the interpretability and control offered by classical iterative schemes.
@@ 1.2 Inverse Problems and Regularization
where $K:\mathcal{X}\rightarrow\mathcal{Y}$ is a compact operator between Hilbert spaces, and $w$ represents measurement noise. The ill-posed nature manifests through three distinct aspects of the operator-theoretic framework: either injectivity or surjectivity of the forward operator may not hold, or stability of the solution map might be violated. For instance, if $K$ is a compact operator with an infinite-dimensional range, then surjectivity and stability are violated. This is, for instance, the case for the ray transform operator that underlies many medical imaging modalities, such as computed tomography (CT) and positron emission tomography (PET). To see this, first consider the case where the range $\mathscr{R}(K)$ of the forward operator $K$ is not closed, implying that solutions may not exist for an arbitrary $y\in\mathcal{Y}$, as any non-zero noise component $w$ orthogonal to $\mathscr{R}(K)$ renders the problem unsolvable in the strict sense. Second, the potential non-triviality of $\mathscr{N}(K)$, the null-space of the forward operator $K$, violates uniqueness, particularly evident in limited-angle tomography where certain features become invisible. Most critically, the unboundedness of the generalized inverse $K^{\dagger}$, when restricted to $\mathscr{R}(K)$, leads to extreme sensitivity to noise, leading to an unstable solution (i.e., a small amount of noise in the measurement yields a drastically different solution).
@@
where the functional $R:\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ encodes prior knowledge about plausible solutions. A solution $x_{\alpha}$ to (3) is alternatively denoted as the output of a parametric reconstruction operator $\mathcal{R}_{\alpha}\coloneqq\mathcal{R}(\cdot,\alpha):\mathcal{Y}\rightarrow\mathcal{X}$. The classical Tikhonov regularization tikhonov_main employs $R(x)=\frac{1}{2}\|x\|_{\mathcal{X}}^{2}$, yielding the explicit solution $x_{\alpha}=(K^{\top}K+\alpha I)^{-1}K^{\top}y$, where $K^{\top}$ denotes the adjoint [^1] We use the notation $K^{\top}$ to denote the adjoint of $K$ regardless of the image domain $\mathcal{X}$. When $\mathcal{X}=\mathbb{R}^{n}$, the adjoint operator $K^{\top}$ reduces to the transpose of the matrix $K$. of the forward operator $K$. This solution corresponds to a spectral filter with coefficients $f_{\alpha}(\sigma)=\sigma/(\sigma^{2}+\alpha)$. While the quadratic ($L^{2}$) penalty on $x$ guarantees stability of the reconstruction, it does so by uniformly penalizing deviations across all frequencies. This effect can be seen from the Euler–Lagrange optimality condition $K^{\top}(Kx-y)+\alpha x=0$, where the regularization parameter $\alpha$ acts as a frequency-independent damping term. The outcome is a systematic suppression of both noise and fine-scale features, resulting in overly smoothed reconstructions and noticeable loss of sharp edge structures. To mitigate this limitation, one may instead consider sparsity-promoting penalties such as the $L^{1}$ norm lasso_tibshirani , typically applied on a transform domain. For instance, to promote sparsity of the wavelet coefficients of the image, one can choose $R(x)=\|Wx\|_{1}$, where $W$ denotes an appropriate wavelet transform operator. Such sparsity-promoting $L^{1}$ norm-based regularizers are non-differentiable, leading to a non-smooth variational optimization problem for reconstruction (which requires iterative solvers, unlike the variational problem with the Tikhonov regularizer admitting a closed-form solution). Another popular and widely adopted choice is the TV regularizer $R(x)=\|\nabla x\|_{1}$ discussed in Section 1.1, which effectively allows the reconstruction to tolerate large local variations and thus preserve sharp discontinuities such as edges. However, the piecewise-constant bias induced by TV regularization tends to replace smooth gradients in the original image by artificial flat regions separated by sharp transitions of intensity RUDIN1992259_tv ; Chambolle1997_tv_related . Moreover, textures and fine oscillatory details tend to be lost, since they are not easily represented in the sparse model enforced on the gradient image by the TV regularizer.
@@ 1.3 Convergent Regularization
To obtain stable solutions to inverse problems, a mechanism is needed to handle varying noise levels in the measurement. When the measurement noise level is large (small), one must apply a stronger (weaker) regularization: this ensures that the variational framework (3) for reconstruction optimally trades off data-fidelity with the prior knowledge through the parameter $\alpha$. The explicit dependence of $\alpha$ on the measurement noise can be explained by interpreting (3) as a Bayesian maximum a-posteriori estimation problem with an image prior proportional to $\exp(-\beta\,R(x))$ and Gaussian measurement noise with variance $\gamma_{w}^{2}$, resulting in $\alpha=\beta\gamma_{w}^{2}$. For this purpose, the concept of convergent regularization has proven highly useful. Regularization can be roughly understood as a convergence requirement to a unique solution, such as the minimum-norm solution $x^{\dagger}$, where convergence occurs as the noise level $\delta\rightarrow 0$. Formally, consider the previously discussed reconstruction operator $\mathcal{R}_{\alpha}\coloneqq\mathcal{R}(\cdot,\alpha)$, which parametrizes a family of continuous operators $\mathcal{R}_{\alpha}:\mathcal{Y}\to\mathcal{X}$. The parameter $\alpha$ depends on the noise level $\delta>0$, where $\|y^{\delta}-y^{0}\|\leq\delta$ and $y^{0}\coloneqq Kx$ denotes noise-free measurement data. We say that the family of reconstruction operators is a convergent regularization method if there exists a parameter choice rule $\alpha=\alpha(\delta,y^{\delta})$ such that reconstructions $x^{\delta}\coloneqq\mathcal{R}_{\alpha(\delta,y^{\delta})}(y^{\delta})$ converge to $x^{\dagger}\coloneqq K^{\dagger}y^{0}$ (given by the pseudo-inverse) as noise vanishes, in the sense that
@@
In other words, we have point-wise convergence of the reconstruction operators to the pseudo-inverse, i.e. $\mathcal{R}_{\alpha(\delta,y^{\delta})}(y^{\delta})\to K^{\dagger}y^{0}$ as $\delta\to 0$. We refer interested readers to engl1996regularization ; scherzer2009variational ; benning2018modern for a detailed discussion on convergent regularization schemes (and several convergence rate results in the classical regularization literature). While this is somewhat restrictive as it only considers convergence to the least-squares minimum-norm solution, this nevertheless serves as an important tool to design learned regularization methods, i.e., learned reconstruction approaches that formally satisfy the above convergence criterion.